# The HST-Hyperion Survey: Environmental Imprints on the Stellar-Mass Function at $z\sim 2.5$

### v1

@@ Appendix A The Structure of Hyperion
In this section, we provide additional details on the structure of Hyperion, shown in Figure 11.
@@ A.1 Peaks of Hyperion
Table 2: Overdensity Peaks in Hyperion – The various overdensity peaks (as defined in Section 3.2) within Hyperion with at least one galaxy in each MC iteration on average. Columns correspond to the (1) peak number, central (2) right ascension, (3) declination, and (4) redshift, (5) total mass, (6) total volume (not corrected for elongation), and (7) average number of galaxies in each MC iteration.
ID $\rm R.A.$ $\rm decl.$ $z$ $M_{\rm tot}$ Volume $\rm\langle N_{gal}\rangle$
@@ A.2 High-Redshift Offshoot
A more likely explanation, however, is that this apparent offshoot is the result of redshift-space “smearing” in the VMC maps. As shown in Figure 9, the fraction of galaxies assigned via photometric redshifts rises from $\sim 40\%$ in the main body ($z\lesssim 2.52$) to $\sim 80\%$ at higher redshift. These galaxies may or may not truly belong to Hyperion, and some likely have broad redshift probability distributions.
@@
Figure 9: Sources of Redshifts in Hyperion – We show, as a function of redshift, the fraction of galaxies within Hyperion across all 100 MC iterations with redshifts determined from photometry (blue), ground-based spectroscopy (orange), and HST-grism spectroscopy (green). The lower-redshift portion of Hyperion is dominated by spectroscopic redshifts, where as the higher-redshift portion is dominated by photometric redshifts.
Figure 10: Voxel Distribution Within Hyperion – We show the distribution of voxels within Hyperion for four different overdensity thresholds: $n_{\sigma}\geq 2.5$ (orange), $n_{\sigma}\geq 3.0$ (yellow), $n_{\sigma}\geq 3.5$ (green) $n_{\sigma}\geq 4.0$ (blue). The low- and high-redshift portions are connected at $z\sim 2.53$ until an overdensity cut of $n_{\sigma}\geq 4.0$.
Figure 11: The Hyperion Proto-supercluster – (Left) A 3D view of Hyperion with the un-MCed galaxies. Each galaxy is color-coded according to if their redshift is determined photometrically (blue), through ground-based spectroscopy (orange), or HST-grism spectroscopy (green). Note that all galaxies are shown here, while only galaxies with $\log_{10}(M_{*}/M_{\odot})\geq 9.5$ are used in the SMFs. Top Right) Hyperion as seen in the sky-plane, with colors again corresponding to redshift source. (Middle Right) The redshift distribution of overdensities for the galaxies within Hyperion, where $n_{\sigma}$ is given in Equation 3. Colors correspond to redshift source. (Bottom Right) The redshift distribution of galaxies in Hyperion.
@@ Appendix B Schechter Fits
Table 3: Schechter Function Fits for Various SMFs – The best-fit Schechter function fits for various SMFs throughout the paper. Parameters for single- and double-Schechter functions correspond to equations 5 and 6, respectively. Columns correspond to: (1) what is being fit, (2) the characteristic mass, (3) power-law slope of single-Schechter fit (4) normalization of single-Schechter fit, (5) power-law slope of double-Schechter fit, (6) normalization for double-Schechter fit, (7) reduced-$\chi^{2}$ value of the fit, (8) Bayesian Information Criterion of the fit (via Equation 7), and (9) the Figure the fit can be seen in.
@@
Single-Schechter $11.13\pm 0.22$ $1.41\pm 0.14$ $0.58\pm 0.34$ $0.807$ $10.28$ 5
Double-Schechter $10.43\pm 0.18$ $1.28\pm 0.25$ $1.67\pm 0.95$ $1.77\pm 1.27$ $0.48\pm 0.58$ $0.505$ $11.91$ 5
Field††††footnotemark: †† $10.97^{+0.06}_{-0.07}$ 1.46-1.46 $0.24^{+0.03}_{-0.02}$ 27.413 5
Field $11.21\pm 0.10$ $1.55\pm 0.05$ $0.12\pm 0.03$ $1.305$ $12.76$ 7
$10.89\pm 0.08$ 1.3-1.3 $0.32\pm 0.04$ $6.271$ $41.79$ 7
Hyperion Outskirts $11.49\pm 0.36$ $1.63\pm 0.10$ $0.15\pm 0.13$ $0.258$ $7.53$ 7
$10.94\pm 0.14$ 1.3-1.3 $0.73\pm 0.14$ $0.671$ $8.18$ 7
Hyperion Peaks $12.00\pm 4.52$ $1.28\pm 0.29$ $0.78\pm 3.36$ $0.784$ $6.40$ 7
$12.37\pm 3.68$ 1.3-1.3 $0.55\pm 1.45$ $0.523$ $4.79$ 7
222
Figure 12: Parameter Spaces of the single-Schechter Fits – To explore the degeneracies of the parameters in the single-Schechter functions, we plot the parameter spaces of the fits to the SMFs of (Columns 1-4) the field, Hyperion as a whole, Hyperion’s outskirts, and Hyperion’s peaks. In each row, we fit over a grid of two of the three parameters in the single-Schechter function, and marginalize over the remaining parameter to minimize the $\chi^{2}$ value. The parameter spaces explore are (Row 1) $\log_{10}M^{*}-\alpha$, (Row 2) $\log_{10}M^{*}-\Phi^{*}$, and (Row 3) $\alpha-\Phi^{*}$. For each subplot, we additionally plot the point with the minimum $\chi^{2}$ value in the given parameter space (magenta X). In each row, we additionally plot a $3\sigma$-contour from the field, corresponding to $\Delta\chi^{2}\simeq 11.8$, to make comparisons between the likelihood surfaces from the field and the other populations in each column easier.
@@ Appendix C SMF Ratio Fits
In order to highlight the difference between the SMF of the peaks in Hyperion and the field, we analyze the log-ratio of the two (see Figure 8). As part of the analysis, we use a simplistic model in which $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*}/M_{\odot})+\log_{10}(A)$. However, given the relative complexity of the Schechter function, it is not immediately obvious why the ratio should take this form. To explore this more, we briefly derive an analytical expression for the ratio, $\log_{10}(n_{\rm gal}^{H}/n_{\rm gal}^{f})$, in terms of the parameters of the two single-Schechter functions. Given Equation 5, and defining $\mu_{H}\equiv\log_{10}(M_{*}/M^{*}_{H})$ and $\mu_{f}\equiv\log_{10}(M_{*}/M^{*}_{f})$, it follows:
$\begin{split}\frac{n_{\rm gal}^{H}}{n_{\rm gal}^{f}}=\frac{\Phi^{*}_{H}}{\Phi^{*}_{f}}\exp\left(10^{-\mu_{H}}+10^{-\mu_{f}}\right)\\ \qquad\qquad\times 10^{\left[(\alpha_{H}+1)\mu_{H}-(\alpha_{f}+1)\mu_{f}\right]}\end{split}$
Upon simplifying this expression, one arrives at an expression for the log-ratio of the SMFs as a function of stellar mass, $M_{*}$, given as:
@@
$+\log_{10}(e)\left[10^{\mu_{f}}-10^{\mu_{H}}\right]+\mathcal{M}_{0}$
where, for brevity, we have introduced the constant:
$\begin{split}\mathcal{M}_{0}\equiv(\alpha_{f}+1)\cdot\log_{10}\left(\frac{M^{*}_{f}}{M_{\odot}}\right)&-(\alpha_{H}+1)\cdot\log_{10}\left(\frac{M^{*}_{H}}{M_{\odot}}\right)\\ &\quad+\log_{10}\left(\frac{\Phi^{*}_{H}}{\Phi^{*}_{f}}\right)\end{split}$
@@
Figure 13: Expected SMF-Ratios – The SMFs of the Hyperion outskirts (magenta) and peaks (orange) after being normalized by the coeval field. The points are identical to those shown in Figure 8. The lines are the expected form of the ratios using the best-fit parameters for the single-Schechter functions. For most the stellar mass range considered in this work, the ratio is expected to be nearly linear in log-log space.

### v2

@@ Appendix A The structure of Hyperion
In this section, we provide additional details on the structure of Hyperion. Figure 10 shows 2D projections of Hyperion in the R.A.-Dec., R.A.-$z$, and Dec.-$z$ planes. Also plotted are the member galaxies of Hyperion (before any Monte Carlo sampling of redshifts) and their redshift sources.
Figure 10: The distribution of Hyperion member galaxies, projected in the (left) R.A.-Dec., (middle) R.A.-$z$, and (right) Dec.-$z$ planes. In each panel, the solid and dashed contours correspond to Hyperion’s outskirts and peaks, respectively, as defined by the VMC overdensity map. Additionally, the locations of the member galaxies are plotted prior to any Monte Carlo redshift sampling, and their marks are distinguished according to the redshift source: COSMOS2020 photometry (blue circles), ground-based spectroscopy (orange squares), or HST/grism spectroscopy (green triangles). For galaxies with both a ground-based spectroscopic redshift and a HST/grism redshift, the redshift with the highest quality according to the quality flags discussed in Section 2 is chosen. The histograms along each axis show the distribution of member galaxies in the corresponding projected dimension, color-coded by redshift source.
@@ A.1 Peaks of Hyperion
Table 2: Overdensity peaks in Hyperion. [^3] The various overdensity peaks (as defined in Section 3.2) within Hyperion with at least one galaxy in each MC iteration on average. The columns correspond to (1) peak number, central; (2) right ascension; (3) declination; (4) redshift; (5) total mass; (6) total volume (not corrected for elongation); and (7) average number of galaxies in each MC iteration.
ID $\rm R.A.$ $\rm Dec.$ $z$ $M_{\rm tot}$ Volume $\rm\langle N_{gal}\rangle$
@@ A.2 High-redshift offshoot
A more likely explanation, however, is that this apparent offshoot is the result of redshift-space smearing in the VMC maps. As shown in Figure 11, the fraction of galaxies assigned via photometric redshifts rises from $\sim 40\%$ in the main body ($z\lesssim 2.52$) to $\sim 80\%$ at higher redshift. These galaxies may or may not truly belong to Hyperion, and some likely have broad redshift probability distributions.
@@
Figure 11: We show, as a function of redshift, the fraction of galaxies within Hyperion across all 100 MC iterations with redshifts determined from photometry (blue), ground-based spectroscopy (orange), and HST-grism spectroscopy (green). The lower-redshift portion of Hyperion is dominated by spectroscopic redshifts, whereas the higher-redshift portion is dominated by photometric redshifts.
Figure 12: We show the distribution of voxels within Hyperion for four different overdensity thresholds: $n_{\sigma}\geq 2.5$ (orange), $n_{\sigma}\geq 3.0$ (yellow), $n_{\sigma}\geq 3.5$ (green), $n_{\sigma}\geq 4.0$ (blue). The low- and high-redshift portions are connected at $z\sim 2.53$ until an overdensity cut of $n_{\sigma}\geq 4.0$.
@@ Appendix B Schechter fits
Table 3: Schechter function fits for various SMFs. [^4] The best-fit Schechter function fits for various SMFs throughout the paper. The parameters for single- and double-Schechter functions correspond to equations 5 and 6, respectively. The clumns correspond to (1) what is being fit, (2) the characteristic mass, (3) the power-law slope of single-Schechter fit, (4) normalization of single-Schechter fit, (5) power-law slope of double-Schechter fit, (6) normalization for double-Schechter fit, (7) reduced-$\chi^{2}$ value of the fit, (8) the Bayesian information criterion of the fit (via Equation 7), and (9) the figure the fit can be seen in.
@@
Single-Schechter $11.13\pm 0.22$ $1.41\pm 0.14$ $0.58\pm 0.34$ $0.807$ $10.28$ 6
Double-Schechter $10.43\pm 0.18$ $1.28\pm 0.25$ $1.67\pm 0.95$ $1.77\pm 1.27$ $0.48\pm 0.58$ $0.505$ $11.91$ 6
Field††††footnotemark: †† $10.97^{+0.06}_{-0.07}$ 1.46-1.46 $0.24^{+0.03}_{-0.02}$ 27.413 6
Field $11.21\pm 0.10$ $1.55\pm 0.05$ $0.12\pm 0.03$ $1.305$ $12.76$ 8
$10.89\pm 0.08$ 1.3-1.3 $0.32\pm 0.04$ $6.271$ $41.79$ 8
Hyperion Outskirts $11.49\pm 0.36$ $1.63\pm 0.10$ $0.15\pm 0.13$ $0.258$ $7.53$ 8
$10.94\pm 0.14$ 1.3-1.3 $0.73\pm 0.14$ $0.671$ $8.18$ 8
Hyperion Peaks $12.00\pm 4.52$ $1.28\pm 0.29$ $0.78\pm 3.36$ $0.784$ $6.40$ 8
$12.37\pm 3.68$ 1.3-1.3 $0.55\pm 1.45$ $0.523$ $4.79$ 8
Figure 13: To explore the degeneracies of the parameters in the single-Schechter functions, we plot the parameter spaces of the fits to the SMFs of (Columns 1-4) the field, Hyperion as a whole, Hyperion’s outskirts, and Hyperion’s peaks. In each row, we fit over a grid of two of the three parameters in the single-Schechter function, and marginalize over the remaining parameter to minimize the $\chi^{2}$ value. The parameter spaces explored are (row 1) $\log_{10}M^{*}-\alpha$, (row 2) $\log_{10}M^{*}-\Phi^{*}$, and (row 3) $\alpha-\Phi^{*}$. For each subplot, we plot the point with the minimum $\chi^{2}$ value in the given parameter space (magenta X). In each row, we additionally plot a $3\sigma$ contour from the field, corresponding to $\Delta\chi^{2}\simeq 11.8$, to make comparisons between the likelihood surfaces from the field and the other populations in each column easier.
@@ Appendix C SMF ratio fits
In order to highlight the difference between the SMF of the peaks in Hyperion and the field, we analyze the log-ratio of the two (see Figure 9). As part of the analysis, we use a simplistic model in which $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*}/M_{\odot})+\log_{10}(A)$. However, given the relative complexity of the Schechter function, it is not immediately obvious why the ratio should take this form. To explore this more, we briefly derive an analytical expression for the ratio, $\log_{10}(n_{\rm gal}^{H}/n_{\rm gal}^{f})$, in terms of the parameters of the two single-Schechter functions. Given Equation 5, and defining $\mu_{H}\equiv\log_{10}(M_{*}/M^{*}_{H})$ and $\mu_{f}\equiv\log_{10}(M_{*}/M^{*}_{f})$, it follows that
$\begin{split}\frac{n_{\rm gal}^{H}}{n_{\rm gal}^{f}}=\frac{\Phi^{*}_{H}}{\Phi^{*}_{f}}\exp\left(10^{-\mu_{H}}+10^{-\mu_{f}}\right)\\ \qquad\qquad\times 10^{\left[(\alpha_{H}+1)\mu_{H}-(\alpha_{f}+1)\mu_{f}\right]}.\end{split}$
Upon simplifying this expression, one arrives at an expression for the log-ratio of the SMFs as a function of stellar mass, $M_{*}$, given as
@@
$+\log_{10}(e)\left[10^{\mu_{f}}-10^{\mu_{H}}\right]+\mathcal{M}_{0},$
where, for brevity, we have introduced the following constant:
$\begin{split}\mathcal{M}_{0}\equiv(\alpha_{f}+1)\cdot\log_{10}\left(\frac{M^{*}_{f}}{M_{\odot}}\right)&-(\alpha_{H}+1)\cdot\log_{10}\left(\frac{M^{*}_{H}}{M_{\odot}}\right)\\ &\quad+\log_{10}\left(\frac{\Phi^{*}_{H}}{\Phi^{*}_{f}}\right).\end{split}$
@@
Figure 14: The SMFs of the Hyperion outskirts (magenta) and peaks (orange) after being normalized by the coeval field. The points are identical to those shown in Figure 9. The lines are the expected forms of the ratios using the best-fit parameters for the single-Schechter functions. For most the stellar mass range considered in this work, the ratio is expected to be nearly linear in log-log space.